given: x12+2x2-2=0

x1+4x22-4=0

Let (x10;x20)=(1;2) be an initial guess to (x1;x2). Compute the first approximation (x1(1);x2(1)) using the Newton method.


1
Expert's answer
2021-11-08T02:02:03-0500

x1²+2x2=2

x1+4x²=4

let x⁰1,=1 and x⁰2=2

=-2(2)+2=-2

hence first appropriation x¹1=-1 and x¹2= 1